Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4, -5)
________________________
What I have tried
Given - m = -3
(x,y) - x=4 y=-5
y=m(x)+b
-5 = -3(4)+b
im lost from there on.

Found 2 solutions by CharlesG2, Earlsdon:
Answer by CharlesG2(834)

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (4, -5)
Equation of line is y=mx+b where m is the slope and b is the y-intercept (or vertical intercept or (0,b)).
y = -3x + b (plug in (4,-5) and solve for b)
-5 = -3(4) + b
-5 = -12 + b
-5 + 12 = b
7 = b
y = -3x + 7

Answer by Earlsdon(6294)

Ok, you've done everything right so far.
-5 = -3(4)+b Simplify the right side of the equation.
-5 = -12+b Now add 12 to both sides.
7 = b Now write the final equation in slope-intercept form: y = mx+b.
y = -3x+7